OpenStudy (elenathehomeschooler):

What is csc (2pi/3)?

OpenStudy (boldjon):

1.15

OpenStudy (jdoe0001):

check your Unit Circle :)

OpenStudy (elenathehomeschooler):

is it -1/2, sqr3/2 ?

OpenStudy (elenathehomeschooler):

@jdoe0001

OpenStudy (danjs):

recall, get the sine value maybe csc(a) = 1 / sin(a)

OpenStudy (elenathehomeschooler):

i dont get it

OpenStudy (danjs):

that is what it is, csc(angle) = 1 / sin(angle) the sine value is the y coordinate for points on this circle, look for angle 2pi/3 and get the value for sine

OpenStudy (elenathehomeschooler):

its 120 degrees

OpenStudy (danjs):

yes, need the value of the coordinate on the circle there, sine is the y value

OpenStudy (danjs):

since the radius is a 1, each point on the circle is (x,y) = (cos(a) , sin(a))

OpenStudy (danjs):

at 2pi/3 angle, sin(2pi/3) = root(3)/2

OpenStudy (danjs):

you see that value there in the point for y

OpenStudy (danjs):

i have to run... you have the sine value now, use it for the csc value csc(2pi/3) = 1 / sin(2pi/3) = 1 / [root(3)/2] = 2/ root(3) rationalize the denominator if you want to 2root(3) / 3